Searches for Physics Beyond the Standard Model

at CDF

CDF Collaboration

presented by Pekka K. Sinervo June 22, 1992

1. Introduction and Overview

2. Brief Desciption of CDF

3. Precision Electroweak Tests

4. Searches for Phenomena Outside the SM

5. Summary

Introduction

Several ways of probing “Beyond the Standard Model”

• Precision EWK tests

• Studies of phenomena sensitive to “new” physics

• Search for phenomena not predicted by the SM

High energy pp̄ collider is a unique laboratory

• produce anything that couples to quarks and gluons

• large parton-parton √ŝ and Q2

• large integrated luminosities

The CDF Experiment

The Collider Detector at Fermilab

• 350 physicists from 26 institutions

• “general-purpose” detector studying 1.8 TeV p p collisions – excellent charged particle detection

– hermitic calorimetry covering |η| < 4.2 – electron, photon and muon identification

Took data in 1988-89

• Accumulated 4.2 pb−1

• ∼ 2 × 1011 p p interactions

• selected and recorded ∼ 6 × 106 events in real-time

CDF Data Sets

CDF data were classified into distinct data sets offline:

• Inclusive electron sample

• Inclusive muon sample

• High PT (> 120 GeV/c) jets

• Events with large missing ET (E̸T > 25 GeV)

• Dimuon events (e.g. 2 μ with PT > 3 GeV/c)

Used standard “cuts” to define canonical event samples

• Physics analyses used these as basis for most studies

• Avoid discussing details...

Precision Standard Model Tests

W boson mass is an important SM test

• Stringent test of the EW radiative corrections

• Unique to pp̄ colliders (Spp̄S and Tevatron)

• Requires careful control of systematic uncertainties

M_{W} inferred from “transverse mass” distribution for lν_{l} events

• ∼ 1100 e±νe candidates and ∼ 600 μνμ candidates

• Dominate systematic uncertainties arise from – structure functions

– underlying event corrections

⇒ M_{W} = 79.91 ± 0.39 GeV/c^{2}

Quark Compositeness

QCD assumes that quarks are point-like particles

• Quark sub-structure can be modelled by introducing a 4fermion interaction with an energy scale Λ

• Would enhance rate of qq̄ → qq̄ scattering at large Q2

Select inclusive jet events and compare P_{T} spectra with perturbative QCD

• Requires detailed understanding of jet energy calibration • CDF relies on combination of

– in situ measurement of calorimeter non-linearity

– extensive test beam studies

– detailed MC studies effects due to jet fragmentation and underlying event

Sets a quark compositeness scale Λ > 1.4 GeV at 95% C.L.

Lepton Substructure

Lepton substructure will modify Drell-Yan production at high Q^{2}.

pp̄ → l^{+} l^{−}

• Model compositeness with the introduction of additional 4fermion contact interaction in Lagrangian

• Predicts excess of l+ l− events at large mass

• Backgrounds expected to be small

Select events with isolated electron pairs

• Z∘ provides excellent calibration signal

• Require lepton PT > 20 GeV/c

Data in good agreement with Drell-Yan prediction

• Sets electron-quark compositeness scale

Λ{LL \atop −} > 2.2 TeV at 95% C.L. Λ{LL \atop +} > 1.7 TeV at 95% C.L.

• Sets muon-quark compositeness scale

Λ{LL \atop −} > 1.6 TeV at 95% C.L. Λ{LL \atop +} > 1.4 TeV at 95% C.L.

SUSY Searches

SUSY remains an attractive extension to the SM

• Provides a unified description of the 3 strongest forces

• Predicts a spectrum of new fermions and bosons

• Spurred by recent analysis of energy-dependence of gauge couplings

q̃’s and g̃’s couple strongly to p p initial state

• large production cross sections

• decay signature is large E̸T accompanied by additional jets

CDF’s most recent search focuses on decay modes

g̃ → qq̄γ̃

q̃ → qg̃ for Mq̃ > Mg̃
q̃ → qγ̃ for Mg̃ > Mq̃

• Require E̸T > 40 GeV to identify LSP

• QCD Backgrounds reduced by requiring ≥ 2 jets – jet ET > 15 GeV – |η| < 3.5

• Results in 93 candidate events

Expect several SM backgrounds

Z → νν̄X ⇒ 23 ± 8
W → τν̄X ⇒ 41 ± 15
W → μν̄X ⇒ 18 ± 6
W → eν̄X ⇒ 9 ± 3
Heavy Quarks ⇒ 4 ± 4

Conclude that data is consistent with SM sources

• Acceptance for q̃ and g̃ pair production varies with SUSY particle mass

• Also depends on SUSY decay modes – Assumed naive decay model to set SUSY mass limit

• Fit the E̸T distribution to SUSY signal and background

Set SUSY mass limits of

Mg̃ > 152 GeV/c2 at 90% C.L.

Mq̃ > 126 GeV/c2 at 90% C.L.

W′ and Z′ Searches

Many extensions to SM that predict additional gauge groups

• Have two additional “degrees of freedom” – mass of gauge bosons

– couplings of bosons to quark and lepton sectors

CDF has searched for such W′ and Z′ bosons qq̄′ → W′ → lν_{l}

qq̄′ → Z′ → l^{+} l^{−}

Used event sample with very loose lepton ID cuts

• Data is quite clean

• Consistent with tail from W and Z decay

Resulting 95% C.L. limits are

M_{W ′} > 520 GeV/c^{2}
M_{Z′} > 412 GeV/c^{2}

Massive Stable Particles

Searched for evidence of massive, stable particles

• Assumed lifetime γτ > 10−8 s

• Search for slow μ-like particle

• Required PT > 25 GeV/c

Observed 2 candidate events

• To be conservative, assumed both are signal

• Set 95% C.L. limits on cross section

Leptoquark Search

Leptoquarks are predicted in various SM extensions

• couple quarks and leptons of a given generation

• would be pair-produced in p p collisions

• decay to a lepton+quark final state

Searched for evidence of leptoquark production

• Searched for events with dileptons and accompanying jets • Required

– 2 electron candidates with PT > 20 GeV/c – 2 jets with ET > 20 GeV

• Removed Z → e+ e− decays (75 < Mee < 105 GeV/c2)

Observe 0 events

• Can set lower limit on leptoquark mass Mlq

Mlq > 113 GeV/c2 at 95% C.L.

Rare Decays

The radiative decay W^{±} → π^{±}γ

• Suppressed in SM (Γ(W → πγ)/Γ(W → eνe) ∼ 3 × 10−8)

• Sensitive to W − γ coupling and QCD effects

Have some sensitivity to such rare decay modes

• Produced ∼ 24, 000 W + X events

• Inclusive photon used to trigger on W ± → π±γ – Select isolated γ candidate events with PT > 25 GeV/c – Require single charged particle jet with ET > 25 GeV

Selection results in 11 events

• 3 events in signal region 71 < M(πγ) < 89 GeV/c2

• Expect at least 1.4 events in signal region from QCD

Set upper limit

\frac{Γ(W ± → π±γ)}{Γ(W ± → e±νe)} < 7.5 × 10−3 at 95% C.L.

Using SM value for BR(W ± → e±νe) = 0.109 (Mtop > MW) BR(W ± → π±γ) < 8.2 × 10−4

High E_{T} Events?

Also systematically examined the highest E_{T} events

• Studied tail of ∑ ET distribution

• Consistent with perturbative QCD jet production

Future Prospects at CDF

CDF is currently beginning a new data run

• Expect 25 pb−1 by summer 1993

• Expect 100 pb−1 by early 1994

⇒ ×25 increase in data sample

Detector has also been significantly upgraded

• Inner Silicon Vertex Detector (SVX)

• Extended μ coverage to |η| < 1.0

• Incremental improvements to calorimetry and DAQ

Improve the probes into physics of the SM and beyond

• Uncertainty on MW → 0.1 GeV/c2

• Observe top quark production if Mtop < ∼ 150 GeV/c2

• Improve sensitivity to very high-mass phenomena – Do not appear to be background-limited – Obtain factors of 5-10 in cross section limits

• Probe the bottom quark sector

Summary

CDF has not seen any hints of NEW physics Searched for

• Supersymmetric particles (q̃ or g̃)

• W ′ and Z′ gauge bosons

• quark or lepton compositeness

• W → πγ

• Unusual (e.g. high ∑ ET) events

Expect to increase data set by ×25 over next year

• Will increase sensitivity of these searches

• Study rarer processes (e.g. W ±W ∓ production)

• Test consistency of SM with top quark search